this is the equations by substitution.
9x-8y=10
y=-x+3
what is the solution?

Answer by mananth(16946)

9 x -8 y = 10 .............1
1 y = 3 -1 x ---------2
/ 1 =
y = 3 -1 x
Substitute y in (1)
9 x-8 (3-x) = 10
9 x -24 + 8 x = 10
9 x + 8 x = 10 24
17 x = 34
/ 17
x= 2
Plug the value of x in (1)
9 x -8 y = 10
9 * 2 -8 y = 10
18 -8 y = 10
-8 y = -8
/ -8
y= 1
m.ananth@hotmail.ca